# Stability of optimal transport on metric measure spaces

### v1

@@ Abstract
We prove a quantitative stability of Kantorovich potentials on metric measure spaces with lower Ricci curvature bound, thereby confirming a recent conjecture of Kitagawa, Letrouit and Mérigot. Our proof, which employs the heat kernel-regularized $c$-transform, does not rely on linear structure or sectional curvature bounds. As a corollary, we get a quantitative stability of optimal transport maps on Alexandrov spaces with lower curvature bound.
Keywords: optimal transport, heat kernel-regularization, Kantorovich potential, stability, metric measure space, curvature-dimension condition, Ricci curvature
@@ 1 Introduction
1.1 Motivation and setting
Optimal transport.
@@ 1.1 Motivation and setting
$\int_{X}c(x,T(x))\,{\mathrm{d}}\rho(x)=\inf_{H_{\sharp}\rho=\mu}\int_{X}c(x,H(x))\,{\mathrm{d}}\rho(x),$ (MP)
The Monge problem prevents mass splitting, sometimes making the problem ill-posed. Its relaxation, called Kantorovich problem, optimizes over joint couplings instead of deterministic maps. A fundamental result known as Kantorovich duality, asserts that the Kantorovich problem is equivalent to the following dual problem:
@@
In the Euclidean space, for $c(x,y)=\frac{1}{2}|x-y|^{2}$, Brenier’s landmark result [5] established that, for absolutely continuous source measures, the optimal transport map is unique and takes the form $T(x)=\nabla u(x)$ for a convex function $u$. This was later extended to Riemannian manifolds by McCann [32], who showed that for $c(x,y)=\frac{1}{2}{\mathrm{d}}^{2}(x,y)$, the optimal transport map is given almost everywhere by $T(x)=\exp_{x}(-\nabla\varphi(x))$ for a Kantorovich potential $\varphi$.
@@
General setting.
In this paper, we confirm the conjecture of Kitagawa– Letrouit–Mérigot for the following spaces:
  • RCD(K,N)(K,N) spaces (metric measure spaces with synthetic Ricci curvature lower bound): we prove quantitative $L^{1}$-stability of Kantorovich potentials with respect to Wasserstein perturbations of the target measure;
  • Alexandrov spaces with (sectional) curvature lower bound: we establish stability estimates for optimal transport maps in terms of the $L^{2}$-norm of the potential difference.
@@
John domains, named after F. John [23], encompass many cases of interest, such as bounded Lipschitz domains, bounded domains satisfying a cone condition and certain fractal domains (see [7] for more discussions).
Definition 1.1 (John domain).
@@
${\mathrm{d}}(\gamma(t),S^{c})\geq\eta t,~~~\forall t\in[0,\ell(\gamma)],$
where $S^{c}$ denotes the complement of $S$.
In [30, Theorem 1.9], it has been shown that when the source measure $\rho$ is the uniform density on some non-John domain $S\subset\mathbb{R}^{n}$, then no quantitative stability estimates of the form $\|\phi_{\mu}-\phi_{\nu}\|_{L^{2}(\rho)}\leq CW_{p}^{\alpha}(\mu,\nu)$ can hold for any $C,\alpha,p>0$. Moreover, examples in [29] indicates that both unboundedness of the density and openness of $S$ may cause instability of optimal transport maps.
In this paper, we always assume that the Kantorovich potential $\phi_{\mu}$ from $\rho$ to $\mu$ satisfies $\mathbb{E}_{\rho}(\phi_{\mu})=\int_{S}\phi_{\mu}\,{\mathrm{d}}\rho=0$. Together with the uniqueness of the optimal transport map, such $\phi_{\mu}$ is unique and thus makes sense to study its stability.
@@ 1.2 Main results
If $(S,{\mathrm{d}},\mathfrak{m})$ is a compact ${\rm RCD}(K,N)$ space, it supports a global $(1,1)$-Poincaré inequality (cf.[18, 42]) and we can remove the John domain assumption:
@@
Let $(X,{\mathrm{d}})$ be an $n$-dimensional Alexandrov space, $\mathfrak{m}$ be the $n$-Hausdorff measure. Under the same assumption for $S,Y$ and $\rho$ as in Theorem 1.2, and if $S$ additionally has finite perimeter, then there exists a constant $C>0$, depending on $k,n,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),{\rm{Per}}(S),S$, such that for any $\mu,\nu\in\mathcal{P}_{2}(Y)$, we have
@@ 1.3 Strategy: heat kernel-regularized $c$-transform
we make use of heat kernel-regularized $c$-transform (Hopf–Cole transform):
@@
$\mathrm{K}_{t}[\varphi]:=\int_{S}\Phi_{t}[\varphi]\,{\mathrm{d}}\rho.$
@@
Organization. This paper is structured as follows. In Section 2, we prove quantitative stability of Kantorovich potentials on $\mathrm{RCD}(K,N)$ spaces. In Section 3, we establish stability of optimal transport maps on Alexandrov spaces. The Appendix A contains technical lemmas about Poincaré inequalities.
Declaration. The authors declare no conflict of interest and that the manuscript has no associated data.
Acknowledgement. This work is supported by the Ministry of Science & Technology of China (2021YFA1000900, 2021YFA1002200), National Natural Science Foundation of China (12201596) and Shandong Provincial Natural Science Foundation (ZR2025QB05). The authors thank Nicola Gigli, Jun Kitagawa, Nan Li, Cyril Letrouit, Quentin Mérigot, Luca Tamanini for helpful discussions and suggestions on the bibliography.
@@ 2 Stability of Kantorovich potentials
In this section, we study the quantitative stability of Kantorovich potentials on RCD(K,N)(K,N) spaces and prove Theorem 1.2.
@@ 2.1 Heat kernel estimate
We begin by recalling the short-time asymptotic behaviour of the heat kernel on ${\rm RCD}(K,N)$ spaces. It was first studied by Varadhan on Riemannian manifolds [44], and is known as Varadhan formula today.
@@
$-t\log$ $\frac{C_{1}(\epsilon)}{\mathfrak{m}\left(B\big(y,\sqrt{\frac{t}{2}}\big)\right)}+\frac{2{\mathrm{d}}^{2}(x,y)}{4+\epsilon}+\frac{C_{2}(\epsilon)}{2}t^{2}$ (2.3)
$\leq$ $-t\log p_{\frac{t}{2}}(x,y)\leq t\log C_{1}(\epsilon)\mathfrak{m}\left(B\big(y,\sqrt{\frac{t}{2}}\big)\right)+\frac{2{\mathrm{d}}^{2}(x,y)}{4-\epsilon}-\frac{C_{2}(\epsilon)}{2}t^{2}.$
@@ 2.2 Heat kernel regularization
$\mathfrak{m}(B(y_{0},\mathop{\rm diam}\nolimits(Y)+i+1))\leq c_{2}\int_{0}^{\mathop{\rm diam}\nolimits(Y)+i+1}\sinh^{N-1}(\sqrt{\frac{-K}{N-1}}s)\,{\mathrm{d}}s.$
Note that $\sinh^{N-1}(\sqrt{\frac{-K}{N-1}}s)\leq\frac{e^{\sqrt{(N-1)|K|}s}}{2^{N-1}}$, then
$\mathfrak{m}(B(y_{0},\mathop{\rm diam}\nolimits(Y)+i+1))\leq c_{1}\int_{0}^{\mathop{\rm diam}\nolimits(Y)+i+1}\frac{e^{\sqrt{(N-1)|K|}s}}{2^{N-1}}\,{\mathrm{d}}s\leq c_{2}e^{\sqrt{(N-1)|K|}i},$ (2.8)
where $c_{1}=\frac{\mathfrak{m}(B(y_{0},\mathop{\rm diam}\nolimits(Y)))}{\int_{0}^{\mathop{\rm diam}\nolimits(Y)}\sinh^{N-1}(\sqrt{\frac{-K}{N-1}}s){\mathrm{d}}s}$, $c_{2}=c_{1}\frac{e^{\sqrt{(N-1)|K|}(\mathop{\rm diam}\nolimits(Y)+1)}}{2^{N-1}\sqrt{\frac{-K}{N-1}}}$.
@@
By the regularization of heat flow [3, Theorem 6.5], we have $H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}}),H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})\cap\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})\cap L^{\infty}(\mathfrak{m})$. Moreover, by heat kernel estimate (2.2),
$H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)\geq$ $\int_{Y}e^{\frac{\psi(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)$ (2.10)
$\geq$ $e^{\frac{\inf_{Y}\psi}{t}}\int_{Y}\frac{1}{C_{1}(1)\mathfrak{m}(B(y,\sqrt{\frac{t}{2}}))}\exp\left(-\frac{2{\mathrm{d}}^{2}(x,y)}{3t}-\frac{C_{2}(1)}{2}t\right)\,{\mathrm{d}}\mathfrak{m}(y)$
$\geq$ $e^{\frac{\inf_{Y}\psi}{t}}e^{-\frac{2D^{2}}{3t}-\frac{C_{2}(1)}{2}t}\frac{\mathfrak{m}(Y)}{\mathfrak{m}(B(y_{0},\mathop{\rm diam}\nolimits(Y)+1))}\frac{1}{C_{1}(1)}>0.$
So $\frac{1}{H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})}\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})\cap L^{\infty}(\mathfrak{m})$ and by chain rule (see [14, Theorem 4.3.3]), we have $\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$. ∎
@@ 2.3 Gradient estimate
Since $f\in L^{2}(\mathfrak{m})\cap L^{\infty}(\mathfrak{m})$, the heat flow regularization yields $H_{s}(f)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$. Moreover, by [22], we have $p_{s}(\cdot,y)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$ for $\mathfrak{m}$-a.e. $y\in X$. For $\varphi\in\big\{\phi\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})\cap L^{\infty}(\mathfrak{m})\cap{\mathrm{D}}(\Delta):\Delta\phi\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})\cap L^{\infty}(\mathfrak{m})\}$, by integration by parts and Fubini theorem, we get
@@
By approximation (cf. [13, §3.2]) and the Riesz representation theorem for Hilbert module (cf.[14, Theorem 3.2.14, Example 3.2.15 ]) we prove the lemma. ∎
@@
${\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)):=\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\left(\big|\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot))\big|^{2}\right)$
@@
$\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot))=\frac{\int_{X}e^{\frac{\psi_{*}(z)}{t}}\nabla_{x}p_{\frac{t}{2}}(x,z)\,{\mathrm{d}}\mathfrak{m}(z)}{\int_{X}e^{\frac{\psi_{*}(z)}{t}}p_{\frac{t}{2}}(x,z)\,{\mathrm{d}}\mathfrak{m}(z)}.$
@@
By Lemma 2.2, Lemma 2.4 and the chain rule for $L^{2}$-normed modules (cf. [14, Theorem 4.3.3]), we have
$\nabla_{x}\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)=\frac{H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\nabla H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}})-H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}})\nabla H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})}{H^{2}_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})}(x)$
@@
$=$ $\int_{X}\big(v(y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\big)\left(\nabla_{x}\log p_{\frac{t}{2}}(x,y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\right)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)$
@@
$\int_{X}\big(v(y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\big)\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)=0.$
@@
${\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot))=\Delta_{x}\log\left(H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)\right)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\Delta_{x}\log p_{\frac{t}{2}}(x,\cdot)\big).$ (2.16)
@@
${\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot))$ (2.18)
@@
Integrating the Li–Yau type estimate (cf. [21, Theorem 1.2])
$|\nabla_{x}\log p_{\frac{t}{2}}(x,y)|^{2}\leq e^{-\frac{Kt}{3}}\frac{\Delta_{x}p_{\frac{t}{2}}(x,y)}{p_{\frac{t}{2}}(x,y)}+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}.$ (2.20)
@@
$J_{t}(x)\leq e^{-\frac{Kt}{3}}\frac{2\Lambda_{\psi}}{t}\bar{J}_{t}^{\frac{1}{2}}(x)+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}\leq\frac{C_{1}(K,N,\Lambda_{\psi}))}{t^{2}}.$ (2.24)
@@
Take a cut-off function $\eta$ satisfying $\eta\equiv 1$ on $B$, $\eta\equiv 0$ on $X\setminus B(x_{0},2r_{0})$, $0\leq\eta\leq 1$, and $|\nabla\eta|\leq\frac{10}{r_{0}}$. Recall that $a_{1}\mathfrak{m}\leq\rho\leq a_{2}\mathfrak{m}$, by Lemma 2.6, we have
@@
For $I_{2}$, by (2.17) and Li–Yau type estimate (2.20), we have

### v2

@@ Abstract
We prove a quantitative stability of Kantorovich potentials on non-smooth metric measure spaces with synthetic lower Ricci curvature bound, thereby confirming a recent conjecture of Kitagawa, Letrouit and Mérigot. Our proof, which employs the heat kernel-regularized $c$-transform, does not rely on linear structure or sectional curvature bounds, is new even in the smooth setting. As a corollary, we get a quantitative stability of optimal transport maps on Alexandrov spaces with lower curvature bound.
Keywords: optimal transport, heat kernel, Kantorovich potential, quantitative stability, metric measure space, curvature-dimension condition, Ricci curvature
@@ 1 Introduction
1.1 Background and Motivation
@@
$\int_{X}c(x,T(x))\,{\mathrm{d}}\rho(x)=\inf_{H_{\sharp}\rho=\mu}\int_{X}c(x,H(x))\,{\mathrm{d}}\rho(x).$ (MP)
The Monge problem is not always well-posed, since it prevents mass splitting. Its relaxation, called Kantorovich problem, optimizes over joint couplings instead of deterministic maps. By Kantorovich duality, the Kantorovich problem is equivalent to the following dual problem:
@@
In the Euclidean space, for $c(x,y)=\frac{1}{2}|x-y|^{2}$, Brenier’s landmark result [5] established that, for absolutely continuous source measures, the optimal transport map is unique and takes the form $T(x)=\nabla u(x)$ for a convex function $u$. This was later extended to Riemannian manifolds by McCann [32], who showed that for $c(x,y)=\frac{1}{2}{\mathrm{d}}^{2}(x,y)$, the optimal transport map is given by $T(x)=\exp_{x}(-\nabla\varphi(x))$ for a Kantorovich potential $\varphi$.
@@
1.2 General Setting
We confirm the conjecture of Kitagawa–Letrouit–Mérigot in the following setting.
A. Metric measure spaces:
@@
B. Source measures:
In [30, Theorem 1.9], it has been shown that when the source measure $\rho$ is the uniform density on some non-John domain $S\subset\mathbb{R}^{n}$, then no quantitative stability estimates of the form
$\|\phi_{\mu}-\phi_{\nu}\|_{L^{2}(\rho)}\leq CW_{p}^{\alpha}(\mu,\nu).$
@@ B. Source measures:
${\mathrm{d}}(\gamma(t),S^{c})\geq\eta t,~~~\forall t\in[0,\ell(\gamma)].$
John domains encompass many cases of interest, such as bounded Lipschitz domains, bounded domains satisfying a cone condition and certain fractal domains (see [7] for more discussions).
[…]
C. Kantorovich potentials:
It is also necessary to (zero-mean) normalize the Kantorovich potential $\phi$ from $\rho$ such that $\mathbb{E}_{\rho}(\phi)=\int_{S}\phi\,{\mathrm{d}}\rho=0$. Together with the uniqueness of the optimal transport map, such $\phi$ is unique and thus makes sense to talk about its stability.
@@ 1.3 Main Results
In particular, if $(S,{\mathrm{d}},\mathfrak{m})$ is a compact ${\rm RCD}(K,N)$ space, $S$ is surely a John domain. So we have the following corollary.
@@
Let $(X,{\mathrm{d}})$ be an $n$-dimensional Alexandrov space with curvature bounded from below by $k$, $\mathfrak{m}$ be the $n$-Hausdorff measure. Under the same assumption for $S,Y$ and $\rho$ as in Theorem 1.1, and if $S$ additionally has finite perimeter, then there exists a constant $C>0$, depending on $k,n,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),S$, such that for any $\mu,\nu\in\mathcal{P}(Y)$,
@@ 1.4 Strategy: heat kernel-regularized $c$-transform
we make use of the following heat kernel-regularized $c$-transform  [^1] We are told by Luca Taminini that in solving the Schrödinger equation with the Cole–Hopf transform, a similar formula will occur.:
@@
$\mathrm{K}_{t}[\psi]:=\int_{S}\Phi_{t}[\psi]\,{\mathrm{d}}\rho.$
@@
Organization. This paper is structured as follows. In Section 2, we prove the quantitative stability of Kantorovich potentials on $\mathrm{RCD}(K,N)$ spaces. In Section 3, we establish the stability of optimal transport maps on Alexandrov spaces. The Appendix A contains technical lemmas about Poincaré inequalities.
Acknowledgement. The authors thank Nicola Gigli, Jun Kitagawa, Nan Li, Cyril Letrouit, Quentin Mérigot, Luca Tamanini for helpful discussions and suggestions on the bibliography.
@@ 2.1 Heat kernel estimate
We begin by recalling the short-time asymptotic behaviour of the heat kernel on $\rm RCD$} spaces. It was first studied by Varadhan on Riemannian manifolds [44], and is known as Varadhan formula today.
@@
$-t\log$ $\frac{C_{1}(\epsilon)}{\mathfrak{m}\left(B\big(y,\sqrt{\frac{t}{2}}\big)\right)}+\frac{2{\mathrm{d}}^{2}(x,y)}{4+\epsilon}-\frac{C_{2}(\epsilon)}{2}t^{2}$ (2.3)
$\leq$ $-t\log p_{\frac{t}{2}}(x,y)\leq t\log C_{1}(\epsilon)\mathfrak{m}\left(B\big(y,\sqrt{\frac{t}{2}}\big)\right)+\frac{2{\mathrm{d}}^{2}(x,y)}{4-\epsilon}+\frac{C_{2}(\epsilon)}{2}t^{2}.$
@@ 2.2 Heat kernel regularization
$\mathfrak{m}\big(B(y_{0},\mathop{\rm diam}\nolimits(Y)+i+1)\big)\leq c_{1}\int_{0}^{\mathop{\rm diam}\nolimits(Y)+i+1}\sinh^{N-1}(\sqrt{\frac{-K}{N-1}}s)\,{\mathrm{d}}s$
where $c_{1}=\frac{\mathfrak{m}(B(y_{0},\mathop{\rm diam}\nolimits(Y)))}{\int_{0}^{\mathop{\rm diam}\nolimits(Y)}\sinh^{N-1}(\sqrt{\frac{-K}{N-1}}s){\mathrm{d}}s}$. Note that $\sinh^{N-1}(\sqrt{\frac{-K}{N-1}}s)\leq\frac{e^{\sqrt{(N-1)|K|}s}}{2^{N-1}}$, then
$\mathfrak{m}\big(B(y_{0},\mathop{\rm diam}\nolimits(Y)+i+1)\big)\leq c_{1}\int_{0}^{\mathop{\rm diam}\nolimits(Y)+i+1}\frac{e^{\sqrt{(N-1)|K|}s}}{2^{N-1}}\,{\mathrm{d}}s\leq c_{2}e^{\sqrt{(N-1)|K|}i}$ (2.8)
where $c_{2}=c_{1}\frac{e^{\sqrt{(N-1)|K|}(\mathop{\rm diam}\nolimits(Y)+1)}}{2^{N-1}\sqrt{{(N-1)}|K|}}$.
@@
By the regularization of heat flow [3, Theorem 6.5], we have $H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}}),H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})\cap\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})\cap L^{\infty}(\mathfrak{m})$. Moreover, by the heat kernel estimate (2.2),
$H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)\geq\int_{Y}e^{\frac{\psi(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)>0.$ (2.10)
So $\frac{1}{H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})}\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})\cap L^{\infty}(\mathfrak{m})$ and by chain rule (see [14, Theorem 4.3.3]), we know the function $x\mapsto\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$. ∎
@@ 2.3 Gradient estimate
For $\varphi\in{\rm TestF}:=\big\{\phi\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})\cap L^{\infty}(\mathfrak{m})\cap{\mathrm{D}}(\Delta):\Delta\phi\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})\cap L^{\infty}(\mathfrak{m})\}$, by integration by parts and Fubini theorem, we get
@@
By density of $\rm TestF$} (cf. [13, §3.2]) and the Riesz representation theorem for Hilbert module (cf.[14, Theorem 3.2.14, Example 3.2.15 ]), we prove the lemma. ∎
@@
${\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big):=\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\left(\big|\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot))\big|^{2}\right)$
@@
$\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)=\frac{\int_{X}e^{\frac{\psi_{*}(z)}{t}}\nabla_{x}p_{\frac{t}{2}}(x,z)\,{\mathrm{d}}\mathfrak{m}(z)}{\int_{X}e^{\frac{\psi_{*}(z)}{t}}p_{\frac{t}{2}}(x,z)\,{\mathrm{d}}\mathfrak{m}(z)}.$
@@
First of all, by [22], $p_{s}(\cdot,y)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$ for $\mathfrak{m}$-a.e. $y\in X$ so all the formulas above are well-defined. By Lemma 2.2, Lemma 2.4 and the chain rule for $L^{2}$-normed modules (cf. [14, Theorem 4.3.3]), we have
$\nabla_{x}\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)=\frac{H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\nabla H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}})-H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}})\nabla H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})}{\big(H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\big)^{2}}(x)$
@@
$=$ $\int_{X}\big(v(y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\big)\left(\nabla_{x}\log p_{\frac{t}{2}}(x,y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\right)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y),$
@@
$\int_{X}\big(v(y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\big)\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)$
$=$ $\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\int_{X}\big(v(y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\big)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)=0.$
@@
${\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)=\Delta_{x}\log\left(H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)\right)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\Delta_{x}\log p_{\frac{t}{2}}(x,\cdot)\big).$ (2.16)
@@
${\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)$ (2.18)
@@
Integrating the following Li–Yau type estimate (cf. [21, Theorem 1.2])
$|\nabla_{x}\log p_{\frac{t}{2}}(x,y)|^{2}\leq e^{-\frac{Kt}{3}}\frac{\Delta_{x}p_{\frac{t}{2}}(x,y)}{p_{\frac{t}{2}}(x,y)}+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}$ (2.20)
@@
$J_{t}(x)\leq e^{-\frac{Kt}{3}}\frac{2\Lambda_{\psi}}{t}\bar{J}_{t}^{\frac{1}{2}}(x)+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}\leq\frac{C_{1}(K,N,\Lambda_{\psi})}{t^{2}}.$ (2.24)
@@
[…]
@@
For $I_{2}$, by (2.17) and the Li–Yau type estimate (2.20), we have